OpenStudy (anonymous):

A tortoise is walking in the desert. It walks for meters at a speed of meters per minute. For how many minutes does it walk?

OpenStudy (anonymous):

I apologize the numbers did not copy, The turtle is walking 62.5 meters at a speed of 5 meters per minute, how many minutes does it walk ?

jimthompson5910 (jim_thompson5910):

use the formula d = r*t where d = distance r = rate (or speed) t = time

jimthompson5910 (jim_thompson5910):

in this case, d = 62.5 r = 5 t = unknown (for now)

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

yw, tell me what you get

OpenStudy (anonymous):

i'm working on it, sorry so slow

OpenStudy (anonymous):

12.5 ?

jimthompson5910 (jim_thompson5910):

no worries

jimthompson5910 (jim_thompson5910):

12.5 minutes is correct

OpenStudy (anonymous):

Thanks for your help !!

jimthompson5910 (jim_thompson5910):

you're welcome